Use the Balmer formula to calculate the first four wavelengths, in nm, of the spectrum corresponding to n = 3, 4, 5, and 6. show the details of your work.
Answer #1

Balmer formula is

\lambda = B \left ( \frac{n^{2}}{n^{2}-2^{2}} \right )   B = 364.50682 nm and n an integer >2

for n = 3    \lambda = 364.50682 *9/5 = 656.112276 nm

n= 4          \lambda = 364.50682 *16/12 = 486.00909 nm

n= 5          \lambda = 364.50682 *25/21 = 433.93669 nm

n= 6          \lambda = 364.50682 *36/32 = 410.07017 nm